OpenStudy (loser66):

A nonzero polynomial f(x) is said to be monic if the leading coefficient is equal to 1. Prove that every monic polynomial of positive degree can be written as a prduct of monic irreducible polynomials. Please help

OpenStudy (freckles):

i wonder if this can be done with induction

OpenStudy (freckles):

\[P_1(x)=x-r_1 \text{ is an irreducible poylnomial }\] \[P_2(x)=x^2-(r_1+r_2)x+r_1r_2=(x-r_1)(x-r_2) \text{ any 2nd } \\ \text{ degree monic polynomial of degree 2 can be factored } \\ \text{ as a product of monic irreducible polynomials }\] Now suppose for some integer k we have \[P_k(x)=x^k-(r_1+r_2+ \cdots +r_k)x^{k-1}\\ +(r_1r_2+r_1r_3+\cdots r_1r_k+\cdots r_{k-1}r_k)x^{k-2} \\ -(r_1r_2r_3 + \cdots r_{k-2}r_{k-1}r_k)x^{k-3} \cdots +r_1r_2 \cdots r_k \\ =(x-r_1)(x-r_2) \cdots (x-r_k)\] then show \[P_{k+1}=(x-r_1) \cdots (x-r_{k+1})\]

OpenStudy (anonymous):

Just as a clarification on the problem, are these polynomials over a field? Like the reals or complex numbers?

OpenStudy (freckles):

I was thinking it has to be the over the complex field

OpenStudy (freckles):

that is the way any polynomial can be factorable

OpenStudy (anonymous):

Yeah, I believe it would need to be over a field.

OpenStudy (loser66):

f(x) can be on Q, R or C

OpenStudy (loser66):

or any field

OpenStudy (freckles):

Well when I think have C I think that includes all subsets of C such as R and so on...

OpenStudy (anonymous):

but not all subsets of C are a field.

OpenStudy (freckles):

that is true

OpenStudy (loser66):

How to define P(x2) like this? why can't it be x^2 + ax +b as a general form?

OpenStudy (anonymous):

I think you have to drop the monic condition if you dont have a field. Since you can always divide by the leading coefficient that way.

OpenStudy (loser66):

One more question: Is it not that any polynomial can be reduce to irreducible polynomials with its roots?? if the coefficient of the given one is 1, so as its factors, am I right?

OpenStudy (anonymous):

If you start with a general 2nd degree polynomial: $$x^2+ax+b,$$then you could say something like, "If its irreducible, we are done, if not it splits into: $$x^2+ax+b=f(x)g(x)$$ where the deg(f),deg(g)<2. Then by the inductive hypothesis f and g are irreducible.

OpenStudy (loser66):

For example, P_5 has at most 5 roots, so that P_5 can be reduced tor (ax -r1)(bx-r2).... and if the leading coefficient of given one is 1, it means a*b*.... =1 iff a=b=c=...=1?

OpenStudy (anonymous):

over an arbitrary field, we aren't always going to be able to factor a polynomial into 1st degree terms with the roots. For example: $$x^2+1$$ is irreducible over Q[x] and R[x]. $$x^2-2$$ is irreducible over Q[x].

OpenStudy (freckles):

does P5 have 5 roots (not just at most) are we not looking at the complex numbers?

OpenStudy (loser66):

at most because sometimes, it has double root. And the max number of roots is 5

OpenStudy (freckles):

i mean P5 can have the zero 0 with multiplicty 2 but that still counts as 2 zeros

OpenStudy (loser66):

But it doesn't matter, just terminology.

OpenStudy (freckles):

but any P5 with coefficient of x^5 as 1 can be written as (x-r1)(x-r_2)...(x-r_5)

OpenStudy (loser66):

if the leading coefficient is not 1, like 2x^2 -4 , can not be written as (1x - something) (1x- somethingelse), and both them have the leading 1.

OpenStudy (freckles):

but your thing defined monic polynomials as having the coefficient of the term with highest exponents as 1 so why are we talking about anything else?

OpenStudy (anonymous):

Should we be writing this proof as if we are in the complex numbers? Because the complex numbers have the property that every polynomial of degree n has exactly n roots (counting multiplicities), which makes this problem easier than it should be. If the problem is, "Prove that every monic polynomial over a field can be written as a product of irreducible monic polynomials", then I dont think we should use the factorization (x-r1)(x-r2)... since we may not have that.

OpenStudy (freckles):

first sentence actually defines it @Loser66

OpenStudy (loser66):

@nerdguy2535 no, no, we have to prove it in any field, not just complex

OpenStudy (loser66):

@frecles we are talking about the monic and the irreducible polys

OpenStudy (freckles):

If we need to prove it every field I don't see how that is possible because x^2+1 cannot be factored over the reals or rationals like nerd said

OpenStudy (anonymous):

x^2+1 is irreducible and monic, which is all thats required.

OpenStudy (freckles):

@Loser66 I know in your first sentence it tells we are only talking about polynomials where the leading term has coefficient one so I don't know why we are talking about 2x^2-4

OpenStudy (loser66):

@freckles So?? if x^2 +1 is irreducible poly in R, and its leading coefficient is 1, then it is good enough, right?

OpenStudy (freckles):

right but why are we talking 2x^2-4

OpenStudy (freckles):

i do x^2+1 is irreducible in the reals and rationals sorry on that one

OpenStudy (loser66):

@freckles I give a counterexample if the coefficient is not 1, hihihihi..

OpenStudy (freckles):

a counterexample to what?

OpenStudy (loser66):

counterexample if the coefficient of the given poly is not 1, we cannot have the coefficients of its factors are 1.

OpenStudy (freckles):

but we are only considering if it is 1

OpenStudy (freckles):

or so i thought

OpenStudy (loser66):

Yes, yes, I got you!!

OpenStudy (anonymous):

What definition are you using for irreducibility?

OpenStudy (anonymous):

Like, a polynomial is irreducible if and only if.... otherwise, its reducible.

OpenStudy (loser66):

Ok, let me type it: We say that a polynomial in K[x] is irreducible if its degree is positive and it cannot be written as a product of 2 polynomials each of strictly smaller (positive) degree That is the definition on my book.

OpenStudy (anonymous):

That way we have something to start with. I think the idea of an inductive proof is great. Its just like how you prove that the integers can be factored into primes. I think we want to mimic that somehow.

OpenStudy (anonymous):

So lets try inducting on the degree of the polynomial.

OpenStudy (anonymous):

Our inductive statement is, "Any monic polynomial of degree n (or less) can be factored into monic irreducible polynomials". Do you think this is fine?

OpenStudy (anonymous):

I guess I'm thinking Strong Induction since I put the (or less) in there.

OpenStudy (anonymous):

For the base case, if n=1, we have a monic polynomial of degree 1, which by your definition is clearly irreducible.

OpenStudy (anonymous):

Now we assume the statement holds for all polynomials of degree n or less, and try to show we can do the same for a monic degree n+1 polynomial. If the monic degree n+1 polynomial is already irreducible, we are done. If not, then by your definition is factors into a product of 2 polynomials with strictly less degree. Since the degree of our starting polynomial was n+1, the degrees of these two factors must be n or less. So we can use our inductive hypothesis here.

OpenStudy (anonymous):

So these two factor can each be written in terms of monic irreducible polynomials, and by putting the two irreudcible monic factorizations together, you get a monic irreducible factorization for the degree n+1 polynomial.

OpenStudy (loser66):

I got your induction proof. Thanks a lot. One more question: So that we assume that the factors are monic? Don't we have to prove they are monic?

OpenStudy (anonymous):

Ah, yeah, that could be a problem >.< How do we get around that?

OpenStudy (anonymous):

In my head i was thinking: $$c_nx^n+c_{n-1}x^{n-1}+\cdots +c_1x+c_0$$$$=c_n\left(x^n+\frac{c_{n-1}}{c_n}x^{n-1}+\cdots +\frac{c_{1}}{c_n}x+\frac{c_{0}}{c_n}\right)$$

OpenStudy (loser66):

Since I have a proposition. It says: Any polynomials in K[x] of positive degree can be written as a product of irreducible polynomials with the proof follows So that I think the problem here is "monic"

OpenStudy (anonymous):

Ah yes, factoring out the leading coefficient will work. You know the product of the leading coefficients in your two factors is 1. So factor them out and multiply them to make them go away.

OpenStudy (loser66):

@freckles Now you know why I give 2x^2-4 there, right?

OpenStudy (freckles):

" A nonzero polynomial f(x) is said to be monic if the leading coefficient is equal to 1. Prove that every monic polynomial of positive degree can be written as a prduct of monic irreducible polynomials. " I'm still having a problem with this. It says prove for every poly of pos deg can be written as a product of monic polys. I was thinking this has to be strickly over the complex field. Because x^2+1 is a monic poly of pos deg but cannot be written as a product of monic polys over any other field but the complex field

OpenStudy (freckles):

but I guess we could also say the same for x-1 which cannot be written as a product of monic polys over any field

OpenStudy (loser66):

Why?? x^2 +1 is an irreducible one,( in R, Q) nothing to do with it. in C , we follow your proof

OpenStudy (freckles):

I think I'm just lost

OpenStudy (loser66):

I think you confuse yourself. :)

OpenStudy (loser66):

In R, Q , x^2 +1 is monic irreducible, it is done.

OpenStudy (freckles):

So we are trying to prove: If the monic poly is reducible, then it can be written as a product of irrducible monic polynomials.

OpenStudy (anonymous):

I think the word "product" might by throwing you off. The polynomial x^2+1 can be considered as a product only involving one term.

OpenStudy (loser66):

in C, x^2+1 can be written as (x+i)(x-i), both them are monic and irreducible. Proof done in 1 step

OpenStudy (freckles):

but it is 1*(x^2+1) and 1 is not a monic polynomial

OpenStudy (loser66):

ye, 1 is polynomial degree 0

OpenStudy (freckles):

Ok I think I get it

OpenStudy (anonymous):

This is just like the problem, prove any natural number can be written as a product of primes. 3 is a "product of primes"

OpenStudy (loser66):

ok, back to my problem , please. How to prove the factors are monic?

OpenStudy (anonymous):

Say they aren't monic:$$f(x)=c_nx^n+c_{n-1}x^{n-1}+\cdots+c_1x+c_0$$ $$g(x)=b_nx^n+b_{n-1}x^{n-1}+\cdots+b_1x+b_0$$but you know that: $$c_n\cdot b_n=1$$ since their product is monic (we started with a monic polynomial).

OpenStudy (loser66):

nope

OpenStudy (anonymous):

"nope" tp cn times bn being 1?

OpenStudy (anonymous):

to*

OpenStudy (loser66):

the degree of f and g are not good.

OpenStudy (anonymous):

aha ahaha, yes, you are right

OpenStudy (anonymous):

arbitrary degrees, my bad my bad

OpenStudy (anonymous):

ugh now I have to retype that lol >.<

OpenStudy (loser66):

hihihi just copy and paste

OpenStudy (anonymous):

so say f and g are like: $$f(x)=c_kx^k+c_{k-1}x^{k-1}+\cdots+c_1x+c_0$$ $$f(x)=b_rx^r+b_{r-1}x^{r-1}+\cdots+b_1x+b_0$$ where k+r=n+1, and k and r are at least 1.

OpenStudy (freckles):

i wonder if the remainder theorem can be used here

OpenStudy (freckles):

like as far as is it irreducible not and also is it monic or not

OpenStudy (anonymous):

We know $$c_k\cdot b_r=1$$ since their product is monic. Factor out the leading coefficients: $$f(x)=c_k\left(x^n+\frac{c_{n-1}}{c_k}x^{n-1}+\cdots+\frac{c_1}{c_k}x+\frac{c_0}{c_k}\right)$$ do the same for g.

OpenStudy (anonymous):

then you will have: $$c_k\cdot \left(\text{monic polynomial}\right)b_r\cdot \left(\text{monic polynomial}\right)$$and you can cancel out the ck and br

OpenStudy (anonymous):

oops, my degree is still wrong in that last post >.< Im sure you know whats going on though.

OpenStudy (loser66):

I got you. Thanks a lot for the help. @nerdguy2535 and @freckles

OpenStudy (freckles):

I think nerd helped you way more. I probably just confused everyone more. :p

OpenStudy (freckles):

Good question loser.

OpenStudy (loser66):

If you are interesting, you are welcome at my problems. hihihihi.... To me, it 's not good at all. I got stuck and not many people here can help. I do appreciate if someone makes my question clear or gives me another perspective on it. It helps me understand the concept deeper.